OpenStudy (anonymous):

Find an equation of the line containing the given pair of points. (-6,0) (0,7). What is the equation of the line?

OpenStudy (anonymous):

Can you help me out

OpenStudy (aripotta):

first, find the slope using the slope formula:\[\frac{ y _{2}-y _{1} }{ x _{2}-x _{1} }\]

OpenStudy (anonymous):

-7/6

OpenStudy (aripotta):

\[\frac{ 7-0 }{0-(-6) }=\frac{ 7 }{ 0+6 }=\frac{ 7 }{ 6 }\]

OpenStudy (aripotta):

it's not negative

OpenStudy (anonymous):

ok

OpenStudy (aripotta):

so plug the slope and one of the points into point-slope form, which is y = y1 = m(x - x1)

OpenStudy (aripotta):

oop sorry y - y1 = m(x - x1)

OpenStudy (anonymous):

I am not that good with that slope

OpenStudy (anonymous):

Is the answer 7/6x -6 for the answer

OpenStudy (aripotta):

m is the slope, which we just found out is 7/6. we'll take the first point, (-6,0) and plug that into the formula. y1 is the y part of the point, and x1 is the x part of the point. so y - y1 = m(x - x1) y - 0 = 7/6(x + 6)

OpenStudy (aripotta):

now from there, we'll put it in y-intercept form. y - 0 = 7/6(x + 6) y = 7/6(x + 6) all we have to do is distribute 7/6, so y = 7/6x + 7